OpenStudy (anonymous):

arctan summation thing.. 1 sec

OpenStudy (anonymous):

\[\sum_{k = 1}^{n} \arctan \left( \frac{ 1 }{ k^{2} + k + 1 } \right)\]

OpenStudy (anonymous):

Hmmm, I recall \(\tan^{-1}(x)\) having a very simple MacLaurin series.

OpenStudy (anonymous):

Dunno if that would help though, gotta thing...

OpenStudy (anonymous):

oh no, i have never worked with those before

OpenStudy (anonymous):

if it will help i can research it

OpenStudy (anonymous):

Basically do you want an equation in terms of \(n\) then without any summation?

OpenStudy (anonymous):

Well is this a math question or a toy question?

OpenStudy (anonymous):

oh yes, i am attempting to evaluate the summation the end result should be a term with an n in it

OpenStudy (anonymous):

class question I mean

OpenStudy (anonymous):

which one is categorized under toy?

OpenStudy (anonymous):

Meaning it's not a homework question and you can use any possible method to solve it.

OpenStudy (anonymous):

its not a class question. i am looking at some gre stuff and trying to work on the type of math i have not encountered before

OpenStudy (anonymous):

gre?

OpenStudy (anonymous):

so any ideas are welcome, i am just unsure where to begin

OpenStudy (anonymous):

gre = general revised exam, but specifically the math subject exam

OpenStudy (anonymous):

Wouldn't you have come across MacLaurin if you're talking a test for graduate students?

OpenStudy (anonymous):

depends if you're a math graduate or not

OpenStudy (anonymous):

MacLaurin series basically will tell us that: \[ \Large \tan^{-1}(x) = \sum_{n=0}^{\infty} (-1)^n \frac{x^{2n+1}}{2n+1} \]

OpenStudy (anonymous):

Though I'm not sure this summation will be any easier to be honest.

OpenStudy (anonymous):

It is a geometric series at least.

OpenStudy (anonymous):

Scratch that, it's an alternating geometric series.

OpenStudy (anonymous):

Hmmmmmmm, actually it might just work...

OpenStudy (anonymous):

@binarymimic You really wan to learn your series to understand this question.

OpenStudy (anonymous):

\[ \large \begin{split} \sum_{k = 1}^{n} \arctan \left( \frac{ 1 }{ k^{2} + k + 1 } \right) &= \sum_{k = 1}^{n} \sum_{i=0}^{\infty}(-1)^i \left( \frac{ 1 }{ k^{2} + k + 1 } \right)^{2i+1}\frac{1}{2i+1} \\ &= \sum_{i=0}^{\infty} \sum_{k = 1}^{n} (-1)^i \left( \frac{ 1 }{ k^{2} + k + 1 } \right)^{2i+1}\frac{1}{2i+1} \\ &= \sum_{i=0}^{\infty}\frac{(-1)^i}{2i+1} \sum_{k = 1}^{n} \left( \frac{ 1 }{ k^{2} + k + 1 } \right)^{2i+1} \\ \end{split} \]

OpenStudy (anonymous):

This is just me messing around.

OpenStudy (anonymous):

But it's sort of a dead end now. There is probably some other way.

OpenStudy (anonymous):

well i know \[\arctan \left( \frac{ 1 }{ k^{2} + k + 1 } \right) = \arctan \left( \frac{ (k+1)- k }{ k(k+1) + 1 } \right)\] if there was some arctan identity that would take the form \[\arctan \left( \frac{ x - y }{ x*y + 1 } \right)\] that could help. i'm trying to locate some sort of arctan identity

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/List_of_trigonometric_identities says \[\arctan \alpha - \arctan \beta = \arctan \left( \frac{ \alpha - \beta }{ 1 + \alpha \beta } \right)\]

OpenStudy (anonymous):

wait a sec..

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

What if you differentiate and then integrate?

OpenStudy (anonymous):

@wio what if alpha = k + 1 and beta = k

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

Then it would be: \[ \tan^{-1}(k+1)-\tan^{-1}(k) \]?

OpenStudy (anonymous):

\[\arctan \left( \frac{ (k + 1) - k }{ 1 + k(k+1) } \right)\] yeah

OpenStudy (anonymous):

then the summation jut telescopes

OpenStudy (anonymous):

just*

OpenStudy (anonymous):

arctan(k + 1) - arctan(k) + arctan(k+2) - arctan(k + 1) + ... + arcan(n+1)+ arctan(n) = -arctan(k) + arctan(n+1) ?

OpenStudy (anonymous):

Leaving you with \[ \tan^{-1}(n+1) - \tan^{-1}(1) \]

OpenStudy (anonymous):

well thats silly. are they expecting people to memorize every inverse tangent identity ?

OpenStudy (anonymous):

or every inverse trig identity

OpenStudy (anonymous):

I don't remember ANY inverse trig identities. I only really use the normal trig ones.

OpenStudy (anonymous):

.. i wonder if anyone actually got that question right on a test. anyway thanks a MILLION for your help. i have to read up on taylor and mclaurin series now